# MasconCube: Fast and Accurate Gravity Modeling with an Explicit Representation

### v2

@@ 1 Introduction
The geodesy of irregularly shaped small bodies, particularly asteroids and comets, presents fundamental challenges that have driven decades of research in gravitational field modeling. As deep space exploration missions increasingly target these celestial objects, from previous missions like NEAR-Shoemaker [1, 2] and Hayabusa [3, 4, 5] to current endeavors such as Hera [6], Psyche [7], and Tianwen-2 [8], accurate characterization of their internal mass distributions has become essential for mission planning, proximity operations, and scientific understanding.
@@
Recent advances in machine learning have introduced GeodesyNets [20, 21], neural networks that learn three-dimensional density distributions as continuous functions. These networks represent the density field as a neural implicit function, allowing for both shape reconstruction and internal structure inference without requiring prior shape knowledge. GeodesyNets demonstrate remarkable capability in capturing complex gravitational fields with relative errors below $1\%$ even near asteroid surfaces. However, the training process can be computationally intensive, particularly for complex heterogeneous models, and requires several data points.
Physics-Informed Neural Networks for Gravity Modeling (PINN-GM) [22, 23], with their latest iteration PINN-GM III [24], represent another machine learning approach that incorporates physical constraints directly into neural network training to learn gravitational potential. These models demonstrate advantages in compactness, noise robustness, and sample efficiency, but remain computationally expensive to train. Although PINN-GM shows promise for gravity modeling, their capability for geodesy reconstruction has been speculated, but never empirically tested.
@@ 2.1 Asteroid Models
We selected a fictitious planetesimal body previously used in the literature [20, 24], along with three asteroids for which high-resolution shape models are publicly available: 101955 Bennu, 433 Eros, and 25143 Itokawa. For consistency with prior studies [20], we employed simplified, low-resolution versions of their surface meshes. To operate in a dimensionless framework, each mesh was scaled to fit within a sphere of radius 0.8. The resulting volumes were then discretized into tetrahedra, with a mass concentration (mascon) assigned to each one, having a dimensionless mass equal to its volume. The mascon models were subsequently normalized by dividing each mass by the total mass of the body, ensuring that the overall mass sums to 1.
To create some fictitious internal heterogeneity, we modified the mass values of selected mascons and subsequently re-normalized the total mass to unity. Bennu was partitioned into three horizontal layers: the outer layers were assigned a higher density, while the central region was given a lower density. For Eros, we studied 3 different versions: a homogeneous one, one partitioned in 2 homogeneous regions, and one partitioned in 3. Itokawa was divided into two regions, following a configuration inspired by its observed geodesy. A second version of Itokawa presents instead a smooth transition between densities, For the fictitious planetesimal, two hollow configurations were created by setting the mass of selected mascons to zero and then normalizing the mascon values so that their sum remained equal to one. In one version, the hollow region is centrally located; in the other, referred to as the decentered model, the hollow region is offset from the center. A third version of planetesimal presents a homogeneous distribution. The resulting heterogeneous mascon models are illustrated in Figure LABEL:fig:meshes, where the tetrahedra are color-coded according to their assigned density.
@@
This limitation is intentional. Training our model against a ground truth that also uses a 3D grid would simplify the task artificially and could lead to biased, overly optimistic results.
To enable a fair qualitative comparison, we also create a cubified version of the ground truth. This process maps the tetrahedral data onto the same 3D grid used by our MasconCube. For each grid point, we first identify the mesh tetrahedron that contains it. We then assign a dimensionless mass to that grid point, calculated specifically so that its implied density matches the density of the enclosing tetrahedron.
@@ 2.2 Training
MasconCubes are trained with a self-supervised learning approach to reconstruct the internal mass distribution of asteroids from external gravity measurements, as illustrated in Figure 1. Unlike traditional geodesy methods, MasconCubes formulates the problem as a direct optimization of mass values in a 3D grid of point masses (mascons). The training code, as well as the code to reproduce all the results in this paper, is provided online [25].
@@ 2.2.2 Observation Model
In our formulation, the mascon masses $\{m_{j}\}$ are normalized such that their total sum is one, i.e., $\sum_{j}m_{j}=1$. This normalization focuses the learning process on the relative distribution of mass rather than its absolute scale, which is assumed to be known or fixed.
@@ 3 Results
In this section, we compare MasconCubes with GeodesyNets, PINN-GM-III, and the cubified ground-truth, created as described in Section 2.1, which serves as our gold standard. PINN-GM-III have been re-implemented following the details provided in the original paper to evaluate them on our ground-truth models. For completeness, we also report the original metrics from the paper, in case our re-implementation deviates from the intended performance. To ensure a comprehensive and fair comparison that highlights the strengths and weaknesses of each method, we evaluate them using several metrics. Since MasconCubes require the asteroid’s shape to be known, we use differential training for GeodesyNets to ensure they can also exploit this information.
@@ 3.1 Qualitative Comparison of the Geodetic Models
Figure LABEL:fig:density_eros3 shows the learned distribution for Eros - Three regions and a plot of the cubified ground truth for comparison. While both the GeodesyNet and the MasconCube learn an internal structure similar to the target one, the density distribution learned by PINN-GM III is completely different, with a concentration of density around the center of mass of the center of mass of the asteroid, and almost zero density everywhere else. This behavior of PINN-GM have been observed for all the asteroids in the dataset, leading us to conclude that, while it remains a valid approach to estimate the gravitational acceleration, it is not suited to study the geodesy of a body.
An important difference between the methods is observed in their ability to model smooth density distributions. Figure LABEL:fig:density_itokawacos compares the internal structure of Itokawa - Smooth as learned by MasconCube and GeodesyNet, demonstrating that MasconCube produces a more accurate representation of the smooth density variations.
@@ 3.2 Stokes Coefficients
Table 1 compares the $\text{MAE}_{7}$ of the normalized Stokes coefficients, as defined in Equation 7, predicted by MasconCube, GeodesyNet, and Cubified ground-truth. A more in detail coefficient-wise comparison for Eros - Two Regions is reported in Figure LABEL:fig:stokes. Overall, the Stokes coefficients analysis reinforces the effectiveness of MasconCube in accurately modeling complex gravitational fields. The smaller MAE values and closer alignment with ground-truth coefficients underscore MasconCube’s capability to capture subtle features of the gravitational potential up to high spherical harmonic degrees. This enhanced precision is critical for applications requiring detailed gravitational modeling, such as spacecraft navigation and geophysical studies of irregularly shaped celestial bodies.
@@
Table 1: Mean Absolute Error (MAE) of normalized Stokes coefficients up to degree 7 for different gravitational models and bodies. MasconCube consistently achieves lower errors compared to GeodesyNet, approaching, and often the accuracy of the cubified ground-truth model, which serves as the benchmark. Lowest value for each body is shown in bold. Results demonstrate MasconCube’s ability to closely replicate gravitational fields across varied shape and density models. PINN-GM-III is not included in the table, as its learned internal structure largely diverges from the ground truth, as shown in Figure LABEL:fig:density_eros3, leading to meaningless coefficients.
@@ 3.3 Acceleration
Table 2 reproduces results reported in previous works, mainly from [24], to provide a broader context. Since these numbers come from evaluations on different datasets and under different experimental setups, they are not directly comparable. The MasconCube results, along with GeodesyNet trained using the differential framework, were obtained in this work following our own training and evaluation protocol on the Eros – Two Regions model. This choice yields a dataset conceptually similar to that of PINN-GM III, as both are Eros-inspired. Nonetheless, the two differ, since their ground‑truth gravity fields are synthetic and were generated using different procedures. Additionally, we trained a reduced-complexity MasconCube variant, called MasconCube-S, using a similar amount of parameters and training data as PINN-GM III, to facilitate a direct comparison.
@@
Table 2: Comparison of machine learning-based gravity models. Values for most models are taken directly from the literature (primarily [24]), where they were evaluated on different datasets and setups; no reimplementation was performed. Values for the underlined models are computed in this work.
@@
Table 3: Evaluation metrics of gravitational acceleration prediction accuracy for different models on the Eros – Two Regions dataset. The metrics include Cosine Distance, Euclidean Norm Distance, and Relative Euclidean Norm Distance, where lower values indicate higher accuracy. The cubified ground truth serves as the most precise reference model, while the uniform density model serves as a baseline.
@@ 3.4 Trajectory Simulation
Table 4: Trajectory simulation accuracy comparison: mean relative position errors for different gravity modeling approaches. Bold values indicate best performance for each asteroid. MasconCube demonstrates superior accuracy, except for uniform asteroids, where the uniform model coincides with the ground-truth.
@@ 3.5 Training Time
GeodesyNet $1.75$ 10 00010\,000 $291.67$
PINN-GM III $1.90$ $8192$ $259.41$
MasconCube $0.396$ $1000$ $6.6$
Table 5: Comparison of training step times and total training durations for different approaches on an NVIDIA RTX 2080 Ti.
The reduced computational requirements make MasconCubes particularly well-suited for scenarios requiring rapid model updates or real-time adaptation to new gravitational observations. While on-board training for spacecraft applications remains computationally challenging due to power and hardware constraints, MasconCubes represent a step toward this capability. The simplified computational structure also makes them amenable to implementation on specialized hardware, which are commonly used in space applications due to their radiation tolerance and power efficiency.
@@ 4.1 Requirement for Prior Shape Knowledge
Figure LABEL:fig:shape_free_training shows the learned mass distribution of a MasconCube when the asteroid’s shape is unknown, with training points sampled at different altitude ranges. When training points are sampled throughout the entire space around the asteroid, including close-to-surface points, the mass distribution, and consequently the shape, of the asteroid is correctly learned. However, as we begin sampling only points farther from the asteroid surface, the MasconCube training converges to spherical shapes in regions where no training points are sampled. This behavior is likely linked to the ill-posed nature of the gravity inversion problem, where multiple mass distributions can produce similar gravitational signatures at distant observation points, leading the optimization to favor simpler, more regular configurations in the absence of constraining data.
The challenges are exacerbated by the explicit representation nature of MasconCubes compared to the implicit representation of GeodesyNets. This fundamental difference causes gradients to flow differently to parameters according to their spatial position, making it impossible for the model to properly learn the asteroid’s shape when training points are located too far from the surface, as demonstrated in Figure LABEL:fig:gradient_flow_comparison.

### v3

@@ 1 Introduction
The geodesy of irregularly shaped small bodies, particularly asteroids and comets, presents fundamental challenges that have driven decades of research in gravitational field modeling. As deep space exploration missions increasingly target these celestial objects, from previous missions like NEAR-Shoemaker [31, 42] and Hayabusa [16, 43, 10] to current endeavors such as Hera [9], Psyche [25], and Tianwen-2 [44], accurate characterization of their internal mass distributions has become essential for mission planning, proximity operations, and scientific understanding. Moreover, precise gravity models are essential for studying the evolution of natural systems, where higher-order harmonics can influence features such as dust ring formation [19].
@@
Recent work addresses gravity field estimation through both observational and computational approaches. Concepts like SmallSat swarm gravimetry [18] propose distributed measurements to infer bulk gravitational parameters, while mascon-based models [37] use discretized polyhedral sources for orbital dynamics analyses. These efforts typically emphasize measurement strategies or forward gravitational computation rather than direct inference of internal structure.
Recent advances in machine learning have introduced various neural network architectures for asteroid gravity modeling. Early approaches explored Extreme Learning Machines (ELMs) [11], which employ single-layer feedforward networks trained without iterative tuning to learn gravitational acceleration around irregular bodies, demonstrating fast training times suitable for on-board computation. Hopfield Neural Networks have been investigated [28, 45] for on-board reconstruction of spherical harmonics coefficients, showing comparable performance to Extended Kalman Filtering approaches for parameter estimation with reduced computational requirements. Concurrently, GeodesyNets [15, 35] have emerged as neural implicit representations that learn three-dimensional density distributions as continuous functions, allowing for both shape reconstruction and internal structure inference without requiring prior shape knowledge, demonstrating remarkable capability in capturing complex gravitational fields with relative errors below $1\%$ even near asteroid surfaces. Physics-Informed Neural Networks for Gravity Modeling (PINN-GM) [21, 20], with their latest iteration PINN-GM III [23], represent another machine learning approach that incorporates physical constraints directly into neural network training to learn gravitational potential. These models demonstrate advantages in compactness, noise robustness, and sample efficiency, but remain computationally expensive to train.
@@ 2.1 Asteroid Models
We selected a fictitious planetesimal body previously used in the literature [15, 23], along with three asteroids for which high-resolution shape models are publicly available: 101955 Bennu, 433 Eros, and 25143 Itokawa. For consistency with prior studies [15], we employed simplified, low-resolution versions of their surface meshes. To operate in a dimensionless framework, each surface model was uniformly scaled so that its outermost vertex lay on a sphere of radius 0.8.
From these surface meshes, the internal volumes were discretized into tetrahedra using the TetGen [39] implementation of the Delaunay refinement algorithm [33, 38]. This volumetric meshing step generated a fully connected 3D grid that filled the interior of each asteroid while preserving its surface topology. A mass concentration (mascon) was assigned to the centroid of each tetrahedron, with a dimensionless mass equal to its volume. The mascon models were subsequently normalized by dividing each mass by the total mass of the body, ensuring that the overall mass sums to 1.
To create some fictitious internal heterogeneity, we modified the mass values of selected mascons and subsequently re-normalized the total mass to unity. Bennu was partitioned into three horizontal layers: the outer layers were assigned a higher density, while the central region was given a lower density. For Eros, we studied 3 different versions: a homogeneous one, one partitioned in 2 homogeneous regions, and one partitioned in 3. Itokawa was divided into two regions, following a configuration inspired by its observed geodesy. A second version of Itokawa presents instead a smooth transition between densities, For the fictitious planetesimal, two hollow configurations were created by setting the mass of selected mascons to zero and then normalizing the mascon values so that their sum remained equal to one. In one version, the hollow region is centrally located; in the other, referred to as the decentered model, the hollow region is offset from the center. A third version of planetesimal presents a homogeneous distribution. The resulting heterogeneous mascon models are illustrated in Figure 9(a), where the tetrahedra are color-coded according to their assigned density.
@@
(a) Density distributions of the ground truth mascon models, with tetrahedron color representing local density. The figure compares several configurations, including regional variations (LABEL:fig:mesh_bennu, LABEL:fig:mesh_eros_2, LABEL:fig:mesh_eros_3, LABEL:fig:mesh_itokawa), hollow structures (LABEL:fig:mesh_planetesimal, LABEL:fig:mesh_planetesimal_decentered), and uniform bodies (LABEL:fig:mesh_eros_uniform, LABEL:fig:mesh_planetesimal_uniform), alongside a model with a smooth density gradient (LABEL:fig:mesh_itokawa_smooth). Each model’s total mass is normalized to 1, and models derived from the same base mesh utilize the same color legend to facilitate comparison.
@@
This limitation is intentional. Training our model against a ground truth that also uses a 3D grid would simplify the task artificially and could lead to biased, overly optimistic results. The key concern is that grid-aligned evaluation would enable perfect spatial correspondence between model predictions and ground truth mass elements, allowing the model to achieve artificially high accuracy rather than developing genuine capability to approximate continuous or irregularly distributed mass fields. By maintaining the physically faithful tetrahedral ground truth derived from real asteroid mesh models, our evaluation ensures that MasconCube must genuinely learn to approximate arbitrarily distributed mass rather than exploiting alignment artifacts. This structural mismatch acts as a stress test: if the model performs well despite the representational incompatibility, it demonstrates robustness that should transfer to operational scenarios where the true internal mass distribution is unknown and certainly not grid-aligned.
To enable a fair qualitative comparison, we also create a cubified version of the ground truth. This process maps the tetrahedral data onto the same 3D grid used by our MasconCube. For each grid point, we first identify the mesh tetrahedron that contains it. We then assign a dimensionless mass to that grid point, calculated specifically so that its implied density matches the density of the enclosing tetrahedron. Importantly, this cubified representation serves exclusively for visualization purposes and is never used for training or quantitative evaluation.
@@ 2.2 Training
MasconCubes are trained with a self-supervised learning approach, a paradigm where supervision signals are derived from the data or physical laws themselves rather than manual annotation, to reconstruct the internal mass distribution of asteroids from external gravity measurements, as illustrated in Figure 9. It is necessary to train a dedicated model for each asteroid, as the gravity field and internal structure are uniquely determined by the individual body’s shape and mass distribution.
Unlike traditional geodesy methods, MasconCubes formulates the problem as a direct optimization of mass values in a 3D grid of point masses (mascons). The training code, as well as the code to reproduce all the results in this paper, is provided online [30].
@@ 2.2.1 Mascon Grid Initialization
While different initialization schemes may affect convergence speed, the current initialization approach was empirically found to generally lead to better convergence behavior compared to alternatives evaluated during preliminary experiments. Furthermore, training with different random seeds consistently results in similar converged models, suggesting that the training process is robust to initialization variability.
@@ 2.2.2 Observation Model
While MasconCubes does not employ a traditional neural network architecture, the model can be understood as a parameter optimization problem where the trainable parameters are the $\hat{N}$ mascon masses.
@@
In our formulation, the mascon masses $\{m_{j}\}$ are normalized such that their total sum is one, i.e., $\sum_{j}m_{j}=1$. This normalization focuses the learning process on the relative distribution of mass rather than its absolute scale, which is assumed to be known or fixed. Architecture details are provided in the appendix A.2.
@@ 3 Results
In this section, we compare MasconCubes with GeodesyNets, PINN-GM-III, and the cubified ground-truth, created as described in Section 2.1, which serves as our gold standard. PINN-GM-III have been re-implemented following the details provided in the original paper to evaluate them on our ground-truth models. For completeness, we also report the original metrics from the paper, in case our re-implementation deviates from the intended performance. To ensure a comprehensive and fair comparison that highlights the strengths and weaknesses of each method, we evaluate them using several metrics.
MasconCubes require the asteroid’s shape model to be known a priori. To ensure fair comparison, we exclusively compare against the differential training version of GeodesyNets, which also exploits shape information to improve accuracy and efficiency. While PINN-GM III cannot exploit a priori knowledge of the shape of studied body, it requires the Stokes coefficient $\tilde{C}_{20}$ as prior knowledge. Thus, all three methods leverage geometric priors. Shape requirement is practically realistic: in asteroid exploration missions, shape models are typically obtained from optical imagery or radar during approach phases, before high-fidelity gravity data becomes available from close-proximity tracking.
@@ 3.1 Qualitative Comparison of the Geodetic Models
Figure 13(a) shows the learned distribution for Eros - Three regions and a plot of the cubified ground truth for comparison. While both the GeodesyNet and the MasconCube learn an internal structure similar to the target one, the density distribution learned by PINN-GM III is completely different, with a concentration of density around the center of mass of the center of mass of the asteroid, and almost zero density everywhere else. This behavior of PINN-GM have been observed for all the asteroids in the dataset, leading us to conclude that, while it remains a valid approach to estimate the gravitational acceleration, it is not suited to study the geodesy of a body.
(a) Internal structure comparison for Eros – Three Regions model across different methods.
An important difference between the methods is observed in their ability to model smooth density distributions. Figure 13(b) compares the internal structure of Itokawa - Smooth as learned by MasconCube and GeodesyNet, demonstrating that MasconCube produces a more accurate representation of the smooth density variations.
(b) Internal structure comparison for Itokawa – Smooth model across different methods.
@@ 3.2 Stokes Coefficients
Table 1 compares the $\text{MAE}_{7}$ of the normalized Stokes coefficients, as defined in Equation 7, predicted by MasconCube, GeodesyNet, and Cubified ground-truth. A more in detail coefficient-wise comparison for Eros - Two Regions is reported in Figure 16(a). Overall, the Stokes coefficients analysis reinforces the effectiveness of MasconCube in accurately modeling complex gravitational fields. The smaller MAE values and closer alignment with ground-truth coefficients underscore MasconCube’s capability to capture subtle features of the gravitational potential up to high spherical harmonic degrees. This enhanced precision is critical for applications requiring detailed gravitational modeling, such as spacecraft navigation and geophysical studies of irregularly shaped celestial bodies.
Table 1: Mean Absolute Error (MAE) of normalized Stokes coefficients up to degree 7 for different gravitational models and bodies. PINN-GM-III is excluded due to its divergent internal structure (see Figure 13(a)). Lowest values per body are in bold.
@@
(a) Comparison of normalized Stokes coefficients for the Eros – Two Regions model. (a) and (b) show visualizations of the Stokes coefficients obtained from MasconCube and GeodesyNet respectively, illustrating the closer alignment of MasconCube coefficients with the ground truth. (c) plots the Mean Absolute Error $\text{MAE}_{n}$ of the coefficients as a function of degree $n$. The plot highlights MasconCube’s consistently lower error across degrees, confirming its superior fidelity in modeling the gravitational potential.
@@ 3.3 Acceleration
Table 2 reproduces results reported in previous works, mainly from [23], to provide a broader context. Since the reported acceleration errors come from evaluations on different datasets and under different experimental setups, they are not directly comparable. The MasconCube results, along with GeodesyNet trained using the differential framework, were obtained in this work following our own training and evaluation protocol on the Eros – Two Regions model. This choice yields a dataset conceptually similar to that of PINN-GM III, as both are Eros-inspired. Nonetheless, the two differ, since their ground‑truth gravity fields are synthetic and were generated using different procedures. Additionally, we trained a reduced-complexity MasconCube variant, called MasconCube-S (gride size equal to 40), using a similar amount of parameters and training data as PINN-GM III, to facilitate a direct comparison.
Table 2: Comparison of machine learning-based gravity models. Most values are taken from the literature (mainly [23]), while underlined models are computed in this work.
@@
The comparison with the cubified ground-truth makes emerge an apparent discrepancy between superior Stokes coefficient fidelity in Table 1 and higher acceleration errors in Table 3. This actually stems from fundamental differences in how these metrics evaluate field fidelity. Stokes coefficients measure the harmonic decomposition of the gravitational potential globally. Conversely, acceleration prediction directly evaluates point-wise field values.
@@
Table 3: Evaluation metrics for gravitational acceleration prediction on the Eros – Two Regions dataset.
@@ 3.4 Trajectory Simulation
Table 4: Trajectory simulation accuracy comparison: mean relative position errors for different gravity modeling approaches. Bold values indicate best performance for each asteroid.
@@ 3.5 Training Time
Table 5: Comparison of training time required to reach Table 2’s accuracy on an NVIDIA RTX 2080 Ti.
Model Training step time Num. steps Total training time Batch size Parameters
@@
GeodesyNet $1.75$ 10 00010\,000 $291.67$ $1000$ 80 80080\,800
PINN-GM III $1.90$ $8192$ $259.41$ $4096$ $2211$
MasconCube $0.396$ $1000$ $6.6$ $1000$ 35 60335\,603
MasconCube-S $0.067$ $400$ $0.45$ $100$ $2160$
Table 6: Comparison of training step time with different settings on an NVIDIA RTX 2080 Ti.
[…]
Grid size
25 50 100 200
Batch size 250 0.109 0.112 0.112 0.111
500 0.218 0.218 0.215 0.211
1000 0.397 0.407 0.396 0.397
2000 0.761 0.750 0.766 0.766
The reduced computational requirements make MasconCubes particularly well-suited for scenarios requiring rapid model updates or real-time adaptation to new gravitational observations.
While on-board training for spacecraft applications remains computationally challenging due to power and hardware constraints, MasconCubes represent a significant step toward this capability. Recent advances in space-grade neural network implementations on radiation-hardened processors demonstrate the feasibility of deploying machine learning algorithms on spacecraft for autonomous operations [24, 8], including on-board gravity inversion [32]. MasconCubes’ simplified computational architecture, requiring only direct point-mass gravity calculations without iterative numerical integration, substantially reduces the gap between current space-based computing capabilities and on-board adaptive gravity modeling. Moreover, the structure of MasconCubes makes them amenable to implementation on specialized hardware accelerators that are commonly deployed in space applications for their radiation tolerance and power efficiency. Such implementations could enable future missions to perform autonomous gravity field refinement during proximity operations around small bodies, advancing the state-of-the-art in spacecraft autonomy.
@@ 4.1 Requirement for Prior Shape Knowledge
Figure 21(a) shows the learned mass distribution of a MasconCube when the asteroid’s shape is unknown, with training points sampled at different altitude ranges. When training points are sampled throughout the entire space around the asteroid, including close-to-surface points, the mass distribution, and consequently the shape, of the asteroid is correctly learned. However, as we begin sampling only points farther from the asteroid surface, the MasconCube training converges to spherical shapes in regions where no training points are sampled. This behavior is likely linked to the ill-posed nature of the gravity inversion problem, where multiple mass distributions can produce similar gravitational signatures at distant observation points, leading the optimization to favor simpler, more regular configurations in the absence of constraining data.
(a) Shape-free MasconCube training results on Eros for different observation altitude ranges. Black points represent mascons with size proportional to mass. (a) Full altitude range [0, 1] correctly recovers asteroid shape. (b) Mid-range altitudes [0.3, 1] show partial degradation. (c) High altitudes only [0.6, 1] converge to spherical distribution. The progressive loss of shape fidelity demonstrates the challenge of gravity inversion from distant observations.
The challenges are exacerbated by the explicit representation nature of MasconCubes compared to the implicit representation of GeodesyNets. This fundamental difference causes gradients to flow differently to parameters according to their spatial position, making it impossible for the model to properly learn the asteroid’s shape when training points are located too far from the surface, as demonstrated in Figure 21(b).
(b) Gradient flow visualization in MasconCube training demonstrating the spatial localization effect. Red dots indicate regions where mascon masses should increase, blue dots indicate regions where masses should decrease, with dot size proportional to gradient magnitude. This spatial localization explains why MasconCubes struggle with shape-free training when observation points are distant from the asteroid surface, as only mascons near training points receive meaningful updates, leading to convergence toward spherical configurations in unobserved regions.